# Kanpur 208016, India

### v1

@@ 1 Introduction
To move away from this configuration one may consider other physical states. DDF [16] provides a way to realize a generic physical state in the string spectrum, using the DDF vertex operator that takes the form
@@ 2.1 Spin coefficients at finite $s$
$c_{2m-1}|_{s\to\infty}=$ ${1\over(m!)^{2}}\alpha mr_{1}^{-m}\left(\alpha\left(\alpha r_{1}-\alpha r_{2}+r_{2}\right)\right){}^{m-1}\left(\left(r_{1}-r_{2}\right)\,_{2}F_{1}\left(1-m,-m;1;{r_{1}-\alpha r_{1}\over\alpha r_{2}}\right)\right.$
@@ 3 String profile
$\begin{gathered}X^{0}=\tau=t,~{}\vec{X}(t,\sigma)=\left(X^{1}(t,\sigma),X^{2}(t,\sigma),X^{3}(t,\sigma)\right)={1\over 2}\left(\vec{F}(t+\sigma)+\vec{F}(t-\sigma)\right),\\ \vec{F}(u)=\left(f_{1}(u),f_{2}(u),f_{3}(u)\right),\vec{F}^{\prime}(u+2\pi a)=\vec{F}^{\prime}(u).\end{gathered}$ (3.1)
with the constraints:
$\dot{\vec{X}}\cdot\vec{X}^{\prime}=0,~{}(\dot{\vec{X}})^{2}+(\vec{X}^{\prime})^{2}=1,~{}\vec{X}^{\prime}|_{\sigma=0}=\vec{X}^{\prime}|_{\sigma=\pi a}=0.$ (3.2)
The string solution $\vec{X}$ expressed in (3.1) satisfies the first constraint in (3.2). Applying the third constraint at $\sigma=\pi a$, one obtains:
$f_{1}^{\prime}(t+\pi a)=f_{1}^{\prime}(t-\pi a),~{}f_{2}^{\prime}(t+\pi a)=f_{2}^{\prime}(t-\pi a),~{}f_{3}^{\prime}(t+\pi a)=f_{3}^{\prime}(t-\pi a).$ (3.3)
Then at $\sigma=\pi a$, the second constraint reads:
$f_{1}^{\prime}(t-\pi a)^{2}+f_{2}^{\prime}(t-\pi a)^{2}+f_{3}^{\prime}(t-\pi a)^{2}=1.$ (3.4)
For kinematics we choose the total 4-momentum of the string: $p_{1}=(E,\mathbf{0})$, the wave vector: $k^{\mu}$, with the constraints,
$k^{2}=\varepsilon_{k}\cdot k=p_{1}\cdot k=0.$ (3.5)
Satisfying these constraints, one can choose;
$k^{\mu}=E_{\gamma}(0,i,0,i,-\sqrt{2}),~{}\varepsilon_{k}^{\mu}={1\over\sqrt{2}}(1,0,-1,0,0).$ (3.6)
The EM current can be expressed as,
@@
with,
$n^{\mu}(x)=(\dot{X^{0}}(t,\pi a),\dot{\vec{X}}(t,\pi a),0)=\left(1,f_{1}^{\prime}(t-\pi a),f_{2}^{\prime}(t-\pi a),f_{3}^{\prime}(t-\pi a),0\right).$ (3.8)
Thus we have,
@@ Appendix A Details of $\langle{\Phi^{s}_{1}\Phi^{s}_{2}\gamma}\rangle$ computation
Each of the DDF operators $A_{-r}$ acting on the $i$-th vacuum state have DDF photon momentum $-r~{}q_{i}$. For the $|\Phi^{s}_{i}\rangle$ states considered in (2) all the DDF photon momenta are parallel i.e. $q_{1}=-q_{2}=q_{3}=q$ with $q^{2}=0$. This implies that:
@@ A.1 OPE of DDFs
The only non-zero contractions between the several constituent elements of the vertex operators are listed down below. We have simplified our notation further using the fact that the $b_{i}(r,n_{i})$-s contain only the Schür polynomials having $-{i~{}r~{}q\cdot\partial^{s}X_{i}\over s!}$ as arguments. This part can only be contracted with the $e^{i~{}p_{j}\cdot X_{j}}$ part of other vertex operators ($i\neq j$) for our restricted kinematic choice, eq.(A.3). These contractions of $b_{i}$’s result only in numerical factors which we have directly written down in what follows. Ultimately, the full contraction simplifies into the following six different possible complete contractions:
@@ A.2 Amplitude Calculation
The other terms are similarly obtained via similar combinatorics. In the $2$nd term the $a_{3}$ is contracted with either of the $a_{i}(r_{1})$-s and in the $3$rd term $a_{3}$ contracted with either of the $a_{i}(r_{2})$-s. Since $a_{i}(r_{1})$ cannot be contracted with $a_{j}(r_{2})$ when $r_{1}\neq r_{2}$, these are the only possible contractions for our kinematic choice.
@@ Appendix B Summation identities
A further change of variables to $n=m-k$ decouples the sums:
@@
Figure 4: Details of the reordering of the summation indices. The red dots denote the $(m,k)$ indices of the summation. From the first to the second diagram, one counts along the $m$ axis instead of the $k$ axis. In the last figure, one counts the different values of $n=m-k$ for a fixed $k$.

### v2

@@ 1 Introduction
To move away from this configuration one may consider other physical states. DDF [16] provides a way to realize a generic physical state in the string spectrum. The very recent resurgence in study of scattering amplitudes involving these states have started since the works [39, 40]. The DDF vertex operator that takes the form
@@ 2.1 Spin coefficients at finite $s$
[…]
@@ 3 String profile
$\begin{gathered}X^{0}=\tau=t,~\vec{X}(t,\sigma)=\left(X^{1}(t,\sigma),X^{2}(t,\sigma),X^{3}(t,\sigma)\right)={1\over 2}\left(\vec{F}(t+\sigma)+\vec{F}(t-\sigma)\right),\\ \vec{F}(u)=\left(f_{1}(u),f_{2}(u),f_{3}(u)\right),\vec{F}^{\prime}(u+2\pi a)=\vec{F}^{\prime}(u),\end{gathered}$ (3.1)
supplemented with the Virasoro constraints, which imply the following relationships among the embedding functions:
$f_{1}^{\prime}(t+\pi a)=f_{1}^{\prime}(t-\pi a),~f_{2}^{\prime}(t+\pi a)=f_{2}^{\prime}(t-\pi a),~f_{3}^{\prime}(t+\pi a)=f_{3}^{\prime}(t-\pi a),$ (3.2)
$f_{1}^{\prime}(t-\pi a)^{2}+f_{2}^{\prime}(t-\pi a)^{2}+f_{3}^{\prime}(t-\pi a)^{2}=1.$
For kinematics we go to the rest frame of the massive string: $p_{1}=(E,\mathbf{0})$. We choose the wave vector and the polarization, with the constraints, $k^{2}=\varepsilon_{k}\cdot k=p_{1}\cdot k=0$ to be $k^{\mu}=E_{\gamma}(0,i,0,i,-\sqrt{2}),~\varepsilon_{k}^{\mu}={1\over\sqrt{2}}(1,0,-1,0,0)$. The EM current can be expressed as,
@@
where, $n^{\mu}(x)=(\dot{X^{0}}(t,\pi a),\dot{\vec{X}}(t,\pi a),0)$. Thus we have,
@@ Appendix A Details of $\langle{\Phi^{s}_{1}\Phi^{s}_{2}\gamma}\rangle$ computation
Each of the DDF operators $A_{-r}$ acting on the $i$-th vacuum state have DDF photon momentum $r~q_{i}$. For the $|\Phi^{s}_{i}\rangle$ states considered in (2) all the DDF photon momenta are parallel i.e. $q_{1}=-q_{2}=q_{3}=q$ with $q^{2}=0$. This implies that:
@@ A.1 OPE of DDFs
The only non-zero contractions between the several constituent elements of the vertex operators are listed down below. We have simplified our notation further using the fact that the $b_{i}(r,n_{i})$-s contain only the Schür polynomials having $i~r~q\cdot\partial^{s}X_{i}\over s!$} as arguments. This part can only be contracted with the $e^{i~p_{j}\cdot X_{j}}$ part of other vertex operators ($i\neq j$) for our restricted kinematic choice, eq.(A.3). These contractions of $b_{i}$’s result only in numerical factors which we have directly written down in what follows. Ultimately, the full contraction simplifies into the following six different possible complete contractions:
@@ A.2 Amplitude Calculation
The other terms are similarly obtained via similar combinatorics. In the 22nd term the $a_{3}$ is contracted with either of the $a_{i}(r_{1})$-s and in the 33rd term $a_{3}$ contracted with either of the $a_{i}(r_{2})$-s. Since $a_{i}(r_{1})$ cannot be contracted with $a_{j}(r_{2})$ when $r_{1}\neq r_{2}$, these are the only possible contractions for our kinematic choice.
@@ Appendix B Summation identities
A further change of variables to n=mkn=m-k decouples the sums:
@@
Figure 4: Details of the reordering of the summation indices. The red dots denote the $(m,k)$ indices of the summation. From the first to the second diagram, one counts along the $m$ axis instead of the $k$ axis. In the last figure, one counts the different values of n=mkn=m-k for a fixed $k$.